In a theater, orchestra tickets cost $4 more a piece than balcony tickets, and each lodge ticket costs $3 more than an orchestra ticket. Maria paid $508 for four lodge and twelve balcony tickets. How much did each ticket cost?
I am so frazzled on this. If anyone could help me out, thank you!

Answer by jorel1380(3719)

orchestra=balcony+4
lodge=orchestra+3=balcony+4+3=balcony+7
4(b+7)+12b=508
4b+28+12b=508
16b=480
b=30
Balcony tickets are $30; orchestra tickets are $34; and lodge tickets are $37..